GlueX PID Meeting, October 29, 2010

From GlueXWiki
Jump to: navigation, search

Location, Time

Friday, October 29, 2010
2:00 pm EDT
JLab: CEBAF Center, Room F326/7
ESNet: 8542553
EVO: direct meeting link
Conference Group: 800-377-8846, participant code: 77438230
Slides: talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2010-10 on the JLab CUE (you have to be a member of the "halld" group). This directory is accessible from the web at .


  1. Announcements
  2. Minutes from the last meeting
  3. Time-of-Flight
    1. Prototype Status
    2. Calibration system
    3. Mechanical interface to support structure: real estate planning
    4. MOU status
  4. Start Counter Status
    1. MOU plans
  5. Target
    1. MOU status
  6. Schedule review: Mark
  7. Action Items


  • FSU: Sascha Ostrovidov
  • JLab: Eugene Chudakov, Mark Ito (chair), Rory Miskimen, Elton Smith, Sascha Somov, Simon Taylor, Beni Zihlmann

Target Refrigerator Capacity

Chris Keith contacted Mark saying he had contacted the company that makes the refrigerator he has identified. Liquid helium should be no problem with this unit.


The MOU between Florida State and JLab has passed through the Concurrence Chain and has been signed by all parties in Tallahassee. It is on its way to JLab for signatures here.


Rory, Elton, and Mark have agreed on a final draft. That has been sent to Glenn Young for review. After that it goes to Allison Lung and then on the to Concurrence Chain.

TOF Prototype Status

Sasha O. gave us a report on progress at FSU.

  • The machine shop is making a frame for the four-paddle prototype configuration.
    • There was an issue on how to hold the two small paddles in the array. It was decided that the short paddles will be between the long ones.
  • At the time of the last meeting there was a problem with half of the events having missing TDC hits.
    • Often, the overflow bit set on these channels. Sasha spoke with Ed Jastrzembski. The overflow bit is only set if there 16 hits in the TDC window, which in this case is 400 ns.
    • Various things were tried:
      1. reducing trigger window to 50 ns
      2. moving signals to other chips
      3. increasing threshold on the discriminator
    • The overflow problem is largely gone now.
    • There does remain a 10% change that a TDC hit is missing in any given event. This has not been understood.

There was a lot of discussion on various things that could be tried. There seems to be some source of high frequency noise in the system that needs to be tracked down.